A car is being driven along a road at 25 m s-1 when the driver suddenly notices that there is a fallen tree blocking the road 65 m ahead. The driver immediately applies the brakes giving the car a constant acceleration of -5 m s-2.  How far in front of the tree does the car come to rest?


1
Expert's answer
2022-03-20T18:50:16-0400

The distance traveled by a car after the brakes was applied

"d=\frac{v_f^2-v_i^2}{2a}=\frac{0^2-25^2}{2\times (-5)}=62.5\:\rm m"

Hence, the car would be at distance

"65-62.5=2.5\:\rm m"

to the tree.